OpenStudy (anonymous):

Factor 1.5x^2-10x-15

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

1.5 is same as 3/2 so, multiply throughout by 2 and you will have 3x^2-20x -30

OpenStudy (anonymous):

(x+1.26134958403)(x−7.928016250697) this is what i got

OpenStudy (anonymous):

It's not 1.5 it's jus number 1 and than the problem stars of with 5x^2

OpenStudy (anonymous):

haha you qre a clone

OpenStudy (skullpatrol):

5x^2-10x-15

OpenStudy (anonymous):

^^ yes

OpenStudy (skullpatrol):

Factor out the greatest common factor first

OpenStudy (skullpatrol):

$$5x^2-10x-15 =5(x^2-2x-3)$$

OpenStudy (anonymous):

so the answer is \[\left( x-3 \right)\]\[\left( 5x+5 \right)\]

OpenStudy (anonymous):

x=3 or x=-1

OpenStudy (skullpatrol):

@nelotsak wrong answer

OpenStudy (anonymous):

why wrong its correct

OpenStudy (skullpatrol):

The question is to factor 5x^2-10x-15

OpenStudy (skullpatrol):

$$5x^2-10x-15 = 5(x^2-2x-3)$$

OpenStudy (skullpatrol):

@Mikegotdamoolah12 can you finish it?

OpenStudy (anonymous):

okay but \[\left( x-3 \right)\]\[\left( 5x+5 \right)\] are still factors

OpenStudy (skullpatrol):

But it is not equal to zero @nelotsak

OpenStudy (anonymous):

i havnt equated yet

OpenStudy (skullpatrol):

So x=3 or x=-1 is wrong

OpenStudy (anonymous):

its true but the ones in the brackets should be correct